# A Spatial-statistical model to analyse historical rutting data

### 08 Jan 2024 v1

@@ Abstract
Pavement rutting poses a significant challenge in flexible pavements, necessitating costly asphalt resurfacing. To address this issue comprehensively, we propose an advanced Bayesian hierarchical framework of latent Gaussian models with spatial components. Our model provides a thorough diagnostic analysis, pinpointing areas exhibiting unexpectedly high rutting rates. Incorporating spatial and random components, and important explanatory variables like annual average daily traffic (traffic intensity), pavement type, rut depth and lane width, our proposed models account for and estimate the influence of these variables on rutting. This approach not only quantifies uncertainties and discerns locations at the highest risk of requiring maintenance, but also uncover spatial dependencies in rutting (millimetre/year). We apply our models to a data set spanning eleven years (2010-2020). Our findings emphasize the systematic unexpected spatial rutting effect, where more of the rutting variability is explained by including spatial components. Pavement type, in conjunction with traffic intensity, is also found to be the primary driver of rutting. Furthermore, the spatial dependencies uncovered reveal road sections experiencing more than 1 millimeter of rutting beyond annual expectations. This leads to a halving of the expected pavement lifespan in these areas. Our study offers valuable insights, presenting maps indicating expected rutting, and identifying locations with accelerated rutting rates, resulting in a reduction in pavement life expectancy of at least 10 years.
@@ 1 Introduction
Several empirical models have been proposed in the past to predict the performance of pavements (Hu et al. 2022, Yao et al. 2023, Yu et al. 2023) and these models typically include rutting. Empirical models usually relate the accumulation of distresses (e.g., rut, roughness, cracks) in the pavement layers to the load repetitions (annual average daily traffic) and other affecting factors (e.g., pavement or asphalt type, weather, width of the road, layer thickness). Empirical models are known to suffer from limited applicability (Gupta et al. 2014), so when the conditions (for example road construction, weather conditions or damage mechanisms) during the data collection differs from the conditions where the model is applied, one cannot expect to get reasonable results. Empirical models developed for pavement wear in cold climates exists (Jacobson and Wågberg 2005), but the challenge of limited predictability remains present. Calculating the expected rutting for sections based on an empirical model is therefore likely to rather large uncertainty.
@@
Spatial models can provide additional explanation by accounting for trends in location. In the case of pavements, it is likely that sections close to each other behave more similar than sections that are far apart. These location components cannot explain the cause of the behaviour but they do have explanatory value and can be utilized to identify ‘hotspots’, predict future rut depths and quantify uncertainties.
@@
Our primary goal is to develop spatial-statistical models that can pinpoint pavement locations exhibiting unexpected and pronounced spatial rutting effects. To achieve this, we employ a hierarchical Bayesian framework comprising latent Gaussian models (LGMs) that encompass both random and spatial components. Latent Gaussian models have been used in various statistical analyses, as demonstrated by their applications in a multitude of fields (e.g., Dey et al. 2000, Chu et al. 2005, Banerjee et al. 2008, Rue et al. 2009b). These models offer a high degree of flexibility and explicitly account for model dependence. They can approximate entire processes defined over continuous domains and provide probabilistic predictions (Lindgren and Rue 2015). In our specific case, the rutting phenomenon on road surfaces persists regardless of whether it is directly observed, with the latent fields serving as the authentic representation of this rutting. A crucial premise underlying spatial models is that observations situated in close proximity to one another in the spatial domain exhibit greater similarity than those positioned further apart. For a broad class of LGMs, the Integrated Nested Laplace Approximation approach (INLA, Rue et al. 2009a) proves to be an apt tool for performing Bayesian inference.
@@
The proposed models are based on historical rutting observations and a selection of explanatory variables known to influence pavement rutting. These variables include pavement type, Average Annual Daily Traffic (AADT), rut depth from the previous year and lane width. The inclusion of these variables is grounded in existing knowledge of factors that significantly influence rutting in road surfaces (Chan et al. 2010, Lang 2012, Thodesen et al. 2012, Ebrahimi et al. 2019, Yan et al. 2020, Zhao et al. 2020, Pan et al. 2021, Li et al. 2023). Rutting is a complex phenomenon, and the depth of rutting observed in the previous year is a key factor in predicting future rutting. Deeper rut depths from the previous year are expected to correlate with increased rutting in the current year, and this relationship is incorporated into the model. Additionally, lane width is considered as a contributing factor to rutting rates. It is logical to assume that narrower roads are more prone to higher rutting rates due to the limited space for motorists to maneuver. However, it’s important to note that the effect of lane width on rutting is contingent on the intensity of traffic. A critical element in our model is the role of Average Annual Daily Traffic (AADT) on the pavement. Rutting, as we argue, is contingent on the presence of AADT or the load imposed by traffic. Without traffic loading, certain forms of pavement deterioration, such as cracks and roughness, may occur, but rutting is improbable. This hypothesis aligns with prior research on the prediction of pavement variability and deterioration, as illustrated in the work of Hao et al. (2009) and Vedvik (2021) (Hao et al. 2009, Vedvik 2021). In this context, we emphasize a linear approximation for data processing, with the constraint that it passes through the origin. Furthermore, our model is designed with a standardized data approach, where the intercept is set to zero by design, ensuring consistency in the methodology.
@@
As proof of concept, our proposed models are applied to eleven years (2010-2020) of pavement rutting data for one road stretch; the European highway route E14, stretching $67.1$ kilometres from Stjørdal, Norway to Storlien on the Swedish border. We identify areas with unexplained or accelerated rutting and determine which explanatory variables are most important to rutting development. A detailed description of the data used and data cleaning process is given in Section 2. Model notation, formulation and evaluation criteria are given in Section 3. Results and discussion are given in 4. Section 5 explains how road authorities can use the model, based on the results.
@@ 2.3.1 Rut depth and rutting
In the case of repavement, rut depth is smaller the proceeding year resulting in large negative values of $r_{i,t}$. Therefore, following Vedvik (2021), large negative rutting values are removed;
@@
i.e., $r_{i,t}$ is kept if it is at least half the rut depth, or assigned NaN (not a number) otherwise. Some missing rutting data may occur due to uncertainty arising from the actual rut depth measurements or the data collection process (Figure 1, graph (b) shows large variability in rutting).
@@ 2.3.4 Lane width
(a) AADT and Pavement type
@@
(c) Lane width
Figure 1: AADT together with pavement type (a), rutting (for years 2019-2020, mm/year — missing in red) (b), lane width (metre) (c).
@@ 2.4 Exploratory analysis of rutting
On average, over all stretches and years, rutting is $1.67$ mm/year; see Figure  2 (a). The density plots,  2 graph (b), show rutting for all segments for each year. Some years having higher rutting levels and are more widely distributed than others. The lines in graph (c) gives the correlation of rutting values with respect to distance (mean across all years, and two randomly selected years; 2011 and 2013) (Bachmaier and Backes 2011). It is computed from the observed rutting and shows how the rutting values are correlated with respect to distance between the corresponding segments. The distance (metre) is shown on the x-axis, and the calculated correlation between rutting values is on the y-axis. The graph shows some spatial dependencies of pavement rutting along the road. These exist for short distances of up to ($\sim$) $300$ metres (3 kilometres).
Refer to caption (a) Observed rutting Refer to caption (b) Density Refer to caption (c) Autocorrelation
Figure 2: Graph (a): Observed rutting for each year (2011-2020), observed mean rutting $1.67$ mm/year (       ), and the observed mean rutting across segments ( ). Graph (b): Rutting density plots, graph. Graph (c) shows how rutting values are correlated with distance. Spatial dependencies in rutting exist up to 3 kilometres.
@@ 3 Latent Gaussian models for Rutting
A Bayesian latent Gaussian model is a three-stage hierarchical model consisting of a likelihood, a latent Gaussian field, and hyperparameters. In this paper, the true rutting observations $\boldsymbol{\mathrm{r}}=(r_{1},...,r_{n})$ at location $i=1,2...,n$ in year $t$ is modelled as a latent Gaussian field, assumed to have a Gaussian prior with zero mean and a small number of hyperparameters ($<15$). We specify six (6) models with different explanatory variables included and with and without spatial and random effects.
@@ 3.1 Latent field: model for true rutting
We denote the observed rutting for segment $i$ in year $t$ as $r_{i,t}$. Let $\eta_{t}(s_{i})$ denote the true rutting at location $s_{i}$. We model the observed rutting as Gaussian with expected value $\eta_{t}(s_{i})$ and variance $\sigma^{2}_{\varepsilon}$. Hence, $\sigma^{2}_{\varepsilon}$ is the uncertainty of the Gaussian observations. The unstructured error terms $\varepsilon_{i,t}\stackrel{{\scriptstyle i.i.d}}{{\sim}}\mathrm{Normal}(0,\sigma^{2}_{\varepsilon})$ are independent and identically distributed (i.i.d). Further, $r_{i,t}$ is conditionally independent, given the latent field $\eta_{t}(s_{i})$.
3.1.1 Likelihood: observed rutting given true rutting
We first model the observed rutting, given the true rutting. This is referred to as the likelihood model in the Bayesian hierarchical model setting. The likelihood of the rutting, $r_{i,t}$, at road segment $i=1,...,n$ in year $t=1,...,T$ is expressed as
$r_{i,t}\ |\ \eta_{t}(s_{i})\sim\mathrm{N}\left(\eta_{t}(s_{i}).\sigma^{2}_{\varepsilon}\right).$ (3)
We only consider models that have an interaction with AADT, and include a random yearly effect. It is known that more traffic gives more rutting, so we include traffic (AADT) in the model as an interaction with each of the explanatory variables. This means that for pavement type, we assume that the rutting is linear for all pavement types (double traffic means double rutting), but that the different pavement types can have different rates of deterioration. From the explanatory analysis (Figure 2), we found that the rutting seems to be on different levels for different years, and this might be explained by different weather conditions. To account for this, we include a random yearly effect in the model.
@@ 3.1.2 Non-spatial rutting model: based on explanatory variables and year effect
3.1.3 Model including spatial effects
There are several variables not included in our model (Equation (4)) that are known to influence rutting. Examples are variables related to weather, bearing capacity and proportion of heavy traffic load. Most of these variables are spatially dependent, and segments close in space are likely to be more similar than segments further a part. This can be accounted for and utilised by including spatial random effects. The spatial random effect is a Gaussian process (GP) along the road and can be interpreted as the part of the residual (the part of the model not explained by the explanatory variables; lane width, rut depth from the previous year and pavement type) that is spatially dependent. Further, it is reasonable that some of the spatial random effect is consistent over years (e.g., bearing capacity, side slope height, surface and base curvature and proportion of heavy traffic) while others vary between years (e.g., heavy rain events or more heavy traffic due to construction work). Therefore, we include both an annually varying spatial process $\xi_{t}(\boldsymbol{s})$ for year $t=1,...,T$, and a common spatial process $\omega(\boldsymbol{s})$, which we refer to as the spatial rutting component (Ingebrigtsen et al. 2015). The spatial rutting component is the spatial contribution to the total rutting that is not explained by the other variables e.g., variables in the non-spatial model in Equation (4). This contribution can either be positive or negative, depending on the other variables and either over- or underestimate the observed rutting. The linear predictor for the model including spatial effects is described as
@@ 3.1.3 Model including spatial effects
Similarly for the annually varying spatial field $\xi_{t}(\boldsymbol{s})$, the parameters to be estimated are $\sigma^{2}_{\xi_{t}}$ and $\rho_{t}$. This can be done following the approach as in Equation 6. Detail description of the Matérn covariance structure is found in (e.g., Blangiardo and Cameletti 2015, Krainski et al. 2018) and the Supplementary Material LABEL:supplementary:modelformula_SUPPLEMENTARY.
@@ 3.1.4 Other models for comparison
We consider simpler non-spatial and spatial models for comparison. These models include some of the known explanatory variables and are presented in Table 2. The most complex spatial model (in Equation 5) we refer to as Model 1, and it’s non-spatial version, Equation (4), as Model 4. These models include three explanatory variables.
@@ 3.2 Prior models
Vague priors are used, and we let the data inform about the parameters to be estimated. The random yearly effect is evident in density plots in Figure 2 (c), which shows the varying yearly effect of rutting. For easier interpretations, the priors for the spatial components $\xi_{t}(s)$ and $\omega(s)$ are set through the spatial range $\rho$ and marginal standard deviation $\sigma_{\omega}(s)$ of the GRFs with Matérn covariance function in Equation (LABEL:eqn:randge_sigma_2). These are acquired jointly following the penalized complexity (PC) framework (Simpson et al. 2017, Fuglstad et al. 2019). The prior for the range is set through the probability $\mathrm{Pr}(\rho<250)=0.15$, indicating that there is a 15 percent probability that the spatial range is less than 250 metres. For the marginal standard deviation, the prior is set through the probability Pr$(\sigma_{\omega(s)}>2.5)=0.05$, indicating that there is a 5 percent probability that the marginal standard deviation for the rutting is over 2.5 meters. These priors are the same for all the spatial fields.

### 14 Feb 2024 v2

@@ Abstract
Pavement rutting poses a significant challenge in flexible pavements, necessitating costly asphalt resurfacing. To address this issue comprehensively, we propose an advanced Bayesian hierarchical framework of latent Gaussian models with spatial components. Our model provides a thorough diagnostic analysis, pinpointing areas exhibiting unexpectedly high rutting rates. Incorporating spatial and random components, and important explanatory variables like annual average daily traffic (traffic intensity), pavement type, rut depth and lane width, our proposed models account for and estimate the influence of these variables on rutting. This approach not only quantifies uncertainties and discerns locations at the highest risk of requiring maintenance, but also uncover spatial dependencies in rutting (millimetre/year). We apply our models to a data set spanning eleven years (2010-2020). Our findings emphasize the systematic unexplained spatial rutting effect, where some of the rutting variability is accounted for by spatial components. Pavement type, in conjunction with traffic intensity, is also found to be the primary driver of rutting. Furthermore, the spatial dependencies uncovered reveal road sections experiencing more than 1 millimeter of rutting beyond annual expectations. This leads to a halving of the expected pavement lifespan in these areas. Our study offers valuable insights, presenting maps indicating expected rutting, and identifying locations with accelerated rutting rates, resulting in a reduction in pavement life expectancy of at least 10 years.
@@ 1 Introduction
Several empirical models have been proposed in the past to predict the performance of pavements (Hu et al. 2022, Yao et al. 2023, Yu et al. 2023) and these models typically include rutting. Empirical models usually relate the accumulation of distresses (e.g., rut, roughness, cracks) in the pavement layers to the load repetitions (annual average daily traffic) and other affecting factors (e.g., pavement or asphalt type, weather, width of the road, layer thickness). Empirical models are known to suffer from limited applicability (Gupta et al. 2014), so when the conditions (for example road construction, weather conditions or damage mechanisms) during the data collection differs from the conditions where the model is applied, one cannot expect to get reasonable results. Empirical models developed for pavement wear in cold climates exists (Jacobson and Wågberg 2005), but the challenge of limited predictability remains present. Calculating the expected rutting for sections based on an empirical model is therefore likely to give large uncertainty.
@@
Spatial models can provide additional explanation by accounting for trends in location, i.e., along the road. In the case of pavements, it is likely that sections close to each other behave more similar than sections that are far apart. These location components cannot explain the cause of the behaviour but they do have explanatory value and can be utilized to identify ‘hotspots’, predict future rut depths and quantify uncertainties.
@@
Our primary goal is to develop spatial-statistical models that can pinpoint pavement locations exhibiting unexpected and pronounced spatial rutting effects. To achieve this, we employ a hierarchical Bayesian framework comprising latent Gaussian models (LGMs) that encompass both random and spatial components (Jourdain and Klein-Paste 2024). Latent Gaussian models have been used in various statistical analyses, as demonstrated by their applications in a multitude of fields (e.g., Dey et al. 2000, Chu et al. 2005, Banerjee et al. 2008, Rue et al. 2009b). These models offer a high degree of flexibility and explicitly account for model dependence. They can approximate entire processes defined over continuous domains and provide probabilistic predictions (Lindgren and Rue 2015). In our specific case, the rutting phenomenon on road surfaces persists regardless of whether it is directly observed, with the latent fields serving as the authentic representation of this rutting. A crucial premise underlying spatial models is that observations situated in close proximity to one another in the spatial domain exhibit greater similarity than those positioned further apart. For a broad class of LGMs, the Integrated Nested Laplace Approximation approach (INLA, Rue et al. 2009a) proves to be an apt tool for performing Bayesian inference.
@@
The proposed models are based on historical rutting observations and a selection of explanatory variables known to influence pavement rutting that is available for the Norwegian stretch of the road. These variables are pavement type, Average Annual Daily Traffic (AADT), rut depth from the previous year and lane width. The inclusion of these variables is grounded in existing knowledge of factors that significantly influence rutting in road surfaces (Chan et al. 2010, Lang 2012, Thodesen et al. 2012, Ebrahimi et al. 2019, Yan et al. 2020, Zhao et al. 2020, Pan et al. 2021, Li et al. 2023). A critical element in our model is the role of Average Annual Daily Traffic (AADT) on the pavement. Rutting, as we argue, is contingent on the presence of AADT or the load imposed by traffic. Without traffic loading, certain forms of pavement deterioration, such as cracks and roughness, may occur, but rutting is improbable. This hypothesis aligns with prior research on the prediction of pavement variability and deterioration, as illustrated in the work of Hao et al. (2009) and Vedvik (2021). Rutting is a complex phenomenon, and the depth of rutting observed in the previous year is a key factor in predicting future rutting (Archilla and Madanat 2000). Deeper rut depths from the previous year are expected to correlate with increased rutting in the current year, and this relationship is incorporated into the model. Additionally, lane width is considered as a contributing factor to rutting rates. It is logical to assume that narrower roads are more prone to higher rutting rates due to the limited space for motorists to maneuver. However, it’s important to note that the effect of lane width on rutting is contingent on the intensity of traffic. Pavement type also plays a crucial role, particularly in colder climates where it can significantly influence rutting resistance. During winter, rising temperatures lead to shorter periods of frozen ground and snow cover on road surfaces, altering the frequency of freeze-thaw cycles. Consequently, this poses challenges for lower-grade binder types like asphalt gravel concrete (Agc), which are more susceptible to rutting under such conditions
@@
As proof of concept, our proposed models are applied to eleven years (2010-2020) of pavement rutting data for one road stretch; the European highway route E14, stretching $67.1$ kilometres from Stjørdal, Norway to Storlien on the Swedish border. We identify areas with unexplained or accelerated rutting and determine which explanatory variables are most important to rutting development. A detailed description of the data used and data cleaning process is given in Section 2. Model notation, formulation and evaluation criteria are given in Section 3. Results and discussion are given in Section 4. Section 5 explains how road authorities can use the model, based on the results.
@@ 2.3.1 Rut depth and rutting
In the case of repavement, rut depth is smaller in the proceeding year resulting in large negative values of $r_{i,t}$. Therefore, following Vedvik (2021), large negative rutting values are removed;
@@
i.e., $r_{i,t}$ is kept if it is at least half the rut depth, or assigned NaN (not a number) otherwise. Some missing rutting data may occur due to uncertainty arising from the actual rut depth measurements or the data collection process (Figure 1, graph (c) shows large variability in rutting).
@@ 2.3.4 Lane width
(a) AADT/Pavement type, and lane width
@@
Figure 1: AADT together with pavement type, and lane width (a), and rutting (for years 2019-2020, mm/year — missing in red) (b).
@@ 2.4 Exploratory analysis of rutting
On average, over all stretches and years, rutting is $1.62$ mm/year; see Figure  2 (a). The density plots,  2 graph (b), show rutting for all segments for each year. Some years having higher rutting levels and are more widely distributed than others. The lines in graphs (c) and (d) give the correlation of rutting values with respect to distance (mean across all years, and two randomly selected years; 2011 and 2012, respectively) (Bachmaier and Backes 2011). It is computed from the observed rutting and shows how the rutting values are correlated with respect to distance between the corresponding segments. The distance (metre) is shown on the x-axis, and the calculated correlation between rutting values is on the y-axis. The graph shows some spatial dependencies of pavement rutting along the road. These exist for short distances of up to ($\sim$) $300$ metres (3 kilometres).
Refer to caption (a) Observed rutting Refer to caption (b) Density Refer to caption (c) Autocorrelation across years Refer to caption (d) Autocorrelation for years 2011-2012
Figure 2: Graph (a): Observed rutting for each year (2011-2020), observed mean rutting $1.62$ mm/year (     ), and the observed mean rutting across segments ( ). Graph (b): Rutting density plots, graph. Graphs (c) and (d) show how rutting values are correlated with distance. Spatial dependencies in rutting exist up to 3 kilometres.
@@ 3 Latent Gaussian models for Rutting
A Bayesian latent Gaussian model is a three-stage hierarchical model consisting of a likelihood, a latent Gaussian field, and hyperparameters. In this paper, the true rutting at location $s_{i}$ in year $t$ is denoted by $\eta_{t}(s_{i})$ and it is modelled as a latent Gaussian field that includes explanatory variables and spatial effects. We specify six (6) models with different explanatory variables included and with and without spatial and random effects.
3.1 Likelihood: observed rutting given true rutting
We denote the observed rutting for segment $i$ in year $t$ as $r_{i,t}$. The likelihood of the rutting, $r_{i,t}$, at road segment $i=1,...,n$ in year $t=1,...,T$ is expressed as
$r_{i,t}\ |\ \eta_{t}(s_{i})\sim\mathrm{N}\left(\eta_{t}(s_{i}),\sigma^{2}_{\varepsilon}\right).$ (3)
Here, $\sigma^{2}_{\varepsilon}$ is the uncertainty of the Gaussian observations. The unstructured error terms $\varepsilon_{i,t}\stackrel{{\scriptstyle i.i.d}}{{\sim}}\mathrm{Normal}(0,\sigma^{2}_{\varepsilon})$ are independent and identically distributed (i.i.d).
@@ 3.2 Latent field: model for true rutting
We model the observed rutting as Gaussian with expected value $\eta_{t}(s_{i})$ and variance $\sigma^{2}_{\varepsilon}$. Further, $r_{i,t}$ is conditionally independent, given the latent field $\eta_{t}(s_{i})$. We only consider models that have an interaction with AADT, and include a random yearly effect. It is known that more traffic gives more rutting, so we include traffic (AADT) in the model as an interaction with each of the explanatory variables. This means that for pavement type, we assume that the rutting is linear for all pavement types (double traffic means double rutting), but that the different pavement types can have different rates of deterioration. From the explanatory analysis (Figure 2), we found that the rutting seems to be on different levels for different years, and this might be explained by different weather conditions. To account for this, we include a random yearly effect in the model.
@@ 3.2.1 Non-spatial rutting model: based on explanatory variables and year effect
3.2.2 Models including spatial effects
There are several variables not included in our model (Equation (4)) that are known to influence rutting. Examples are variables related to weather, bearing capacity and proportion of heavy traffic load. Most of these variables are spatially dependent, and segments close in space are likely to be more similar than segments further a part. This can be accounted for and utilised by including spatial random effects. The spatial random effect is a Gaussian process (GP) along the road and can be interpreted as the part of the model not explained by the explanatory variables; lane width, rut depth from the previous year and pavement type that is spatially dependent. Further, it is reasonable that some of the spatial random effect is consistent over years (e.g., bearing capacity, side slope height, surface and base curvature and proportion of heavy traffic) while others vary between years (e.g., heavy rain events or more heavy traffic due to construction work). Therefore, we include both an annually varying spatial process $\xi_{t}(\boldsymbol{s})$ for year $t=1,...,T$, and a common spatial process $\omega(\boldsymbol{s})$, which we refer to as the spatial rutting component. The spatial rutting component is the spatial contribution to the total rutting that is not explained by the other variables e.g., variables in the non-spatial model in Equation (4). This contribution can either be positive or negative. The linear predictor for the model including spatial effects is described as
@@ 3.2.2 Models including spatial effects
Similarly for the annually varying spatial field $\xi_{t}(\boldsymbol{s})$, the parameters to be estimated are $\sigma^{2}_{\xi_{t}}$ and $\rho_{t}$. Detail description of the Matérn covariance structure is found in (e.g., Blangiardo and Cameletti 2015, Krainski et al. 2018) and the Supplementary Material LABEL:supplementary:modelformula_SUPPLEMENTARY.
@@ 3.2.3 Other models for comparison
We consider simpler non-spatial and spatial models for comparison. These models include some of the known explanatory variables and are presented in Table 2. The most complex spatial model (in Equation 5) we refer to as Model 1, and it’s non-spatial version, Equation (4), as Model 4. These models include all three explanatory variables.
@@ 3.3 Prior models
Vague priors are used, and we let the data inform about the parameters to be estimated. The random yearly effect is evident in density plots in Figure 2 (c), which shows the varying yearly effect of rutting. For easier interpretations, the priors for the spatial components $\xi_{t}(s)$ and $\omega(s)$ are set through the spatial range $\rho$ and marginal standard deviation $\sigma_{\omega}(s)$ of the GRFs with Matérn covariance function in Equation (LABEL:eqn:randge_sigma_2). These are acquired jointly following the penalized complexity (PC) framework (Simpson et al. 2017, Fuglstad et al. 2019). The prior for the range is set through the probability $\mathrm{Pr}(\rho<250)=0.15$, indicating that there is a 15 percent probability that the spatial range is less than 250 metres. For the marginal standard deviation, the prior is set through the probability Pr$(\sigma_{\omega(s)}>2.5)=0.05$, indicating that there is a 5 percent probability that the marginal standard deviation for the rutting is over 2.5 millimetres. These priors are the same for all the spatial fields.